A compound containing only C, H, and O, was extracted from the bark of the sassafras tree. The combustion of 56.5 mg produced 153 mg of CO2 and 31.4 mg of H2O. The molar mass of the compound was 162 g/mol. Determine its empirical and molecular formulas.
Answer #1

Given

Mass of sample = 56.5 mg = 0.0565 g

Mass of CO 2 = 153 mg = 0.153 g

Mass of H2O = 31.4 mg = 0.0314 g

one molecule of CO 2 contain one C atom.

No. of moles of C = no. of moles of CO 2

We have, no. of moles = Mass / molar mass

No. of moles of C = no. of moles of CO 2 = 0.153 g / 44.010 g /mol

No. of moles of C = no. of moles of CO 2 = 0.003476 mol

Mass of C in the sample = No. of moles of C Molar Mass

Mass of C in the sample = 0.003476 mol   12.01 g / mol = 0.04175 g

One molecule of water contain 2 H atoms.

moles of H = 2 no. of moles of water

  moles of H = 2 no. of moles of water = 2 (0.0314 g / 18.015 g / mol )

moles of H = 0.003486 mol

Mass of H in the sample = moles of H molar mass

Mass of H in the sample = 0.003486 mol 1.0079 g /mol = 0.003514 g

Total mass of C & H in the sample = 0.04175 g + 0.003514 g = 0.04526 g

We have given, Mass of sample = 0.0565 g

Mass of O = Mass of sample - Total mass of C & H in the sample

Mass of O = 0.0565 g - 0.04526 g = 0.01124 g

Moles of O = 0.01124 g / 16.00 g / mol = 0.0007025 mol

Now, calculate ratio of number of moles of C:H:O

0.003476 / 0.0007025 = 4.95 mol C 5

0.003486 / 0.0007025 = 4.96 mol H 5

0.0007025 / 0.0007025 = 1.00 mol O.

In compound, C,H & O are present in the ratio 5 : 5: 1.

Empirical formula of compound is C5H5O.

Now, we can determine molecular formula as shown below.

Empirical formula mass = ( 5 12.01 ) + ( 5 1.0079 ) + ( 1 16.00 ) = 81.09 g / mol

We have relation , Molecular formula = A empirical formula

Where A is the ratio of molar mass of substance to the empirical formula mass.

Therefore,A = Molecular formula mass / empirical formula mass

A = 162.00 g / mol / ( 81.09 g / mol )

A = 1.9977 2

Hence, Molecular formula =A x empirical formula

Molecular formula of compound =2 C5H5O .= C10H10O2 .

ANSWER : Molecular formula of compound is C10H10O2 .